The orbital time of Neptune around the Sun is 165 Earth years and its orbital path is 30 times longer than that of the Earth.


Calculate the orbital speed of Neptune relative to the Earth. orbital speed


State how the speed of travel of Neptune compares with that of the Earth.


1
Expert's answer
2022-06-01T12:49:58-0400
"v=\frac{l}{T}"

"\frac{v_N}{v_E}=\frac{l_N}{l_E}\frac{T_E}{T_N}=\frac{30}{1}*\frac{1}{165}=0.18"

The orbital speed of Neptune is 1/5.5 of orbital speed of Earth.